Find the power of engine that can fire water jet of density 1000 kg/m3
at the
speed of 30m/s upward from the pipe of radius 1cm?
1
Expert's answer
2021-07-30T04:50:04-0400

"P=Q\rho(gh+0.5v^2)\\P=\pi r^2 v\rho(0.5v^2)\\
P=\pi (0.01)^2 (30)(1000)(0.5(30)^2)=4241\ Pa\\"